The second angle of a triangle is 3 times as large as the first. The third angle is 55° more than the first. Find the measure of the smallest angle.

Answer by checkley77(12844)

X+3X+X+55=180
5X=180-55
5X=125
X=125/5
X=25 DEGREES FOR THE FIRST ANGLE.
3*25=75 DEGREES FOR THE SECOND ANGLE.
25+55=80 DEGREES FOR THE THIRD ANGLE.
PROOF:
25+75=80=180
180=180